# Scalable UAV Multi-Hop Networking via Multi-Agent Reinforcement Learning with Large Language Models

### v1

@@ IV-B MARL Framework and Algorithm Overview
In this paper, the proposed model is designed based on the MARL algorithm, where multiple agents interact within a shared environment and iteratively refine their behaviors through trial-and-error. Within the model, each UAV $u$ operates based on its own policy $\pi^{u}(a_{t}^{u}|o_{t}^{u};\theta^{u})$, which dictates the action it should take given its observation $o_{t}^{u}$ and learnable parameters of the policy network $\theta^{u}$. The objective is to find a set of UAV policies that maximizes the joint discounted cumulative reward $\sum_{k=0}^{T-t}\gamma^{k}R_{t+k}$ under the joint policy $\pi(\boldsymbol{a}_{t}|s_{t})=\prod_{u\in\boldsymbol{\mathcal{U}}}\pi^{u}(a_{t}^{u}|o_{t}^{u})$, where $\boldsymbol{a}_{t}$ represents the joint action of all UAVs at time $t$. To train each UAV with a decentralized policy $\pi^{u}$, the training approach is designed based on the Proximal Policy Optimization (PPO) algorithm [40, 41]. The policy objective for each UAV $u$ is initially defined as
@@
where $\zeta_{t}^{\theta^{u}}=\pi^{u}(a_{t}^{u}|o_{t}^{u};\theta^{u})/\pi^{u}(a_{t}^{u}|o_{t}^{u};\theta^{u}_{\text{old}})$ is the probability ratio between the updated and old policies, $A_{t}^{u}$ is the advantage estimation, $\epsilon$ is a hyperparameter that controls the clipping range, restricting policy updates to ensure stable training. The PPO entropy loss and critic loss [40] are also incorporated into the training process. The entropy loss encourages exploration by penalizing deterministic behavior and promoting action diversity. The critic loss is designed to improve value estimation accuracy, providing a more reliable baseline for policy updates and enhancing training stability.
To address the challenges associated with UAV networking, several key modules are proposed based on the MARL framework. To enhance inter-agent cooperation and improve algorithm scalability, an information aggregation module (see Section IV-C) and a grouping strategy (see Section IV-D) are introduced as core components. Meanwhile, the reward mechanism is adjusted to deliver clearer feedback tailored to each agent’s training process (see Section IV-D). During MARL training, the loss function is further augmented with a behavioral constraint term $\mathcal{L}^{\text{BC}}$ to encourage the maintenance of robust multi-hop connectivity (see Section IV-E). Additionally, a knowledge distillation loss $\mathcal{L}^{\text{KD}}$ based on LLMs is designed to promote structured exploration within the MARL paradigm (see Section IV-F). Consequently, the overall objective for each agent $u$ is expressed as
@@ IV-C Information Aggregation for Coordination
To enhance inter-agent cooperation, an information aggregation module is designed for each agent, serving as the input to the policy network. Effective cooperation among UAVs in a multi-agent networking environment relies on the exchange of local observations, enabling more informed decision-making and improved coordination. In the networking problem, UAVs are assumed to be able to share their local observations through the established communication network, enabling each UAV to acquire a more comprehensive view of the environment. The shared information at time $t$ is considered as
[…]
@@
where $\text{concat}(\cdot)$ denotes the concatenation operation, with all vectors flattened before being combined. The local observation of UAV $u$ includes its position $\boldsymbol{l}_{u}$ and the link qualities ${\rho_{u,i}}$ of all links associated with it. The aggregation of observations enables agents to have a more comprehensive understanding of the environment, promoting better coordination and decision-making.
@@ IV-D2 Relay responsibility
where $\text{len}(\cdot)$ calculates the number of communication hops in a path. Additionally, $\boldsymbol{\mathcal{V}}=\{v|v\in\boldsymbol{\mathcal{U}}\backslash\{u\},\rho_{u,v}(t)\geq\rho_{\text{th}},\text{path}_{v}(t)\neq\emptyset\}$ represents the set of UAVs that UAV $u$ can directly connect to, where each UAV in this set can establish a multi-hop backhaul link to the BSs. Formula (22) ensures that UAVs select the link with the fewest hops for data relaying. Based on all communication paths in the multi-hop network, the set of UAVs relayed through UAV $v$ can be defined by
$\boldsymbol{\mathcal{U}}^{\text{RE}}_{u}(t)=\{v|v\in\boldsymbol{\mathcal{U}}\backslash\{u\},u\in\text{path}_{v}(t)\}.$ (23)
@@ IV-E Behavioral Constraint for Robustness
$\boldsymbol{z}_{u}^{\text{BC}}(t)=\frac{\boldsymbol{l}_{g^{*}}(t)-\boldsymbol{l}_{u}(t)}{||\boldsymbol{l}_{g^{*}}(t)-\boldsymbol{l}_{u}(t)||_{2}}.$ (27)
@@
where $\text{map}^{-1}(\cdot)$ represents the inverse of the mapping function $\text{map}(\cdot)$, and $w_{\text{BC}}=||\boldsymbol{l}_{g^{*}}(t)-\boldsymbol{l}_{u}(t)||_{2}$ scales the loss according to the UAV-BS distance. The imposed constraint on UAV behavior ensures stable connections with BSs, preventing erratic movements that may lead to large-scale network disconnections. This regulation is necessary for UAVs directly connected to BSs, as their movement patterns are more deterministic. In contrast, other UAV groups place greater emphasis on non-myopic planning, requiring increased flexibility to adapt to dynamic network conditions, making strict constraints unnecessary.
@@ IV-F LLM Agent and Knowledge Distillation
Given that LLMs possess knowledge and capabilities that are aligned with human preferences, they can be leveraged to guide and enhance the RL training process, thereby reducing unproductive exploration. Pretrained LLMs, although primarily trained on natural language data, can understand high-level semantic information about tasks and scenarios. By embedding chain-of-thought (CoT) reasoning into the inference process, LLM-based agents can interpret tasks, extract key features of networking tasks and environmental states, and make informed decisions. Additionally, random initialization in MARL often leads to inefficient exploration in the early stages, heightening the risk of converging to a local optimum. Given the scarcity of feasible network configurations within the vast search space, we propose utilizing LLM-driven decision-making to enhance the MARL training process. Therefore, an LLM agent is designed to offer an initial attempt at tackling the UAV multi-hop networking task, as shown in Figure 3. To ensure that the LLM fully comprehends the networking task and the environmental state for better reasoning, the networking scenario is simplified to preclude the analysis of thousands of precise numerical values, such as location coordinates. And LLM agents generally do not perform at the level of models that have undergone domain-specific training. In this paper, the LLM module is designed primarily to supply RL agents with decisions that align with common sense, thereby aiding their exploration and training processes. Thus, the adoption of an appropriately simplified environment is considered both acceptable and beneficial. Notably, a single UAV can provide communication services over a contiguous region. In line with this regional characteristic of network service delivery, the entire environment is partitioned into grid cells. The grid width is configured to ensure that the UAV positioned at the center of each grid cell can communicate with the UAVs in the adjacent cells and establish connections with the majority of UEs within the grid, i.e. meeting the SNR requirements. Thus, the grid side length is given by $d^{\text{grid}}=\frac{1}{\sqrt{2}}(\max\{d_{u,v}|\rho_{u,v}\geq\rho_{\text{th}}\})$, where $u,v\in\boldsymbol{\mathcal{U}}$ and the SNR $\rho_{u,v}$ is determined by the distance $d_{u,v}$ between UAVs. Consequently, the distribution of UEs is quantified by counting the number of UEs in each grid cell. Moreover, the UAV positions generated by the LLM are constrained to the centers of these grid cells, reducing the complexity of considering connectivity constraints.
@@
Directly mapping current states to UAV positions poses a challenge for the LLM agent [42, 43], as it requires accounting for UE distribution, organization of multi-hop networks and ensuring UAV connectivity constraints for networking. Inspired by the CoT approach [42], the LLM reasoning process for the problem is partitioned into three sequential steps: a) Analyze UE distribution. Identify densely populated areas and determine candidate UAV deployment locations. b) Address connectivity gaps. Evaluate potential UAV connectivity gaps and reorganize UAV placements to ensure robust network connectivity. c) Determine the final UAV deployment based on the analysis of connectivity constraints and UE distribution. This stepwise approach enables the LLM to use more intermediate tokens to sequentially analyze the scenario and requirements, thereby enhancing the overall quality of the response. Additionally, a verifier is designed to validate the LLM-generated UAV deployment, ensuring it forms a multi-hop network without disconnections. It checks whether the deployment meets predefined thresholds for UAV disconnection and UE connection ratios, adopting it only if the criteria are satisfied. By integrating structured reasoning with validation, this approach enhances the reliability of the LLM output under varying conditions.
@@
$\mathcal{L}^{\text{MATCH}}(\boldsymbol{l}_{u},\boldsymbol{l}^{\text{LLM}}_{\sigma(u)})=||\boldsymbol{l}_{u}-\boldsymbol{l}^{\text{LLM}}_{\sigma(u)}||_{2},$ (31)
@@
$\sigma^{*}=\mathop{\arg\min}_{\sigma\in\mathfrak{S}_{U}}\;\sum_{u\in\boldsymbol{\mathcal{U}}}\mathcal{L}^{\text{MATCH}}(\boldsymbol{l}_{u},\boldsymbol{l}^{\text{LLM}}_{\sigma(u)}).$ (32)
@@
This formulation enables MARL agents to effectively leverage the LLM’s inferred actions as supervisory signals, allowing them to learn the LLM’s decision-making capabilities. This technique promotes structured exploration and enhances the overall training process of MARL.
@@ V-A1 Environment setup
To assess the performance of our proposed method, we employ a simulation environment that spans an area of approximately 3.5 km $\times$ 3.5 km, within which around 150 UEs and 18 UAVs are deployed. The distribution of UEs follows either a uniform pattern or a two-dimensional Gaussian mixture with multiple centers, and their motion follows a Brownian process at a constant velocity. The BSs are strategically positioned at three of the four corners of the area to preclude direct connections with the UEs, thereby simulating communication conditions typically encountered in disaster scenarios. The UAV speed $\omega$ is limited to a maximum displacement of 30 meters within each time slot.

### v2

@@ IV-B MARL Framework and Algorithm Overview
In this paper, the proposed model is designed based on the MARL algorithm, where multiple agents interact within a shared environment and iteratively refine their behaviors through trial-and-error. Within the model, each UAV $u$ operates based on its own policy $\pi^{u}(a_{t}^{u}|o_{t}^{u};\theta^{u})$, which dictates the action it should take given its observation $o_{t}^{u}$ and learnable parameters of the policy network $\theta^{u}$. The objective is to find a set of UAV policies that maximizes the joint discounted cumulative reward $\sum_{k=0}^{T-t}\gamma^{k}R_{t+k}$ under the joint policy $\pi(\boldsymbol{a}_{t}|s_{t})=\prod_{u\in\boldsymbol{\mathcal{U}}}\pi^{u}(a_{t}^{u}|o_{t}^{u})$, where $\boldsymbol{a}_{t}$ represents the joint action of all UAVs at time $t$. To train each UAV with a decentralized policy $\pi^{u}$, the training approach is designed based on the Proximal Policy Optimization (PPO) algorithm [28]. The policy objective for each UAV $u$ is initially defined as
@@
where $\zeta_{t}^{\theta^{u}}=\pi^{u}(a_{t}^{u}|o_{t}^{u};\theta^{u})/\pi^{u}(a_{t}^{u}|o_{t}^{u};\theta^{u}_{\text{old}})$ is the probability ratio between the updated and old policies, $A_{t}^{u}$ is the advantage estimation, $\epsilon$ is a hyperparameter that controls the clipping range, restricting policy updates to ensure stable training. The PPO entropy loss and critic loss [28] are also incorporated into the training process. The entropy loss encourages exploration by penalizing deterministic behavior and promoting action diversity. The critic loss is designed to improve value estimation accuracy, providing a more reliable baseline for policy updates and enhancing training stability. In this work, PPO is adopted due to its stable training behavior, where the clipped policy-update constraint prevents abrupt policy shifts that could otherwise induce disruptive topology changes or unexpected UE disconnections. Building upon PPO, the Independent PPO (IPPO) [10] formulation is adopted as the base framework because its decentralized policy and critic design allow each UAV to make decisions independently from local observations. This structure reduces the coordination complexity that arises when controlling large UAV swarms, mitigates the curse of dimensionality associated with multi-agent state and action spaces, and supports efficient parallel learning. As a result, IPPO offers natural scalability and robustness in dynamic multi-UAV multi-hop networking environments, making it well-suited for the proposed MRLMN architecture.
To address the challenges associated with UAV networking, several key modules are proposed based on the IPPO framework. To enhance inter-agent cooperation and improve algorithm scalability, an information aggregation module (see Section IV-C) and a grouping strategy (see Section IV-D) are introduced as core components. Meanwhile, the reward mechanism is adjusted to deliver clearer feedback tailored to each agent’s training process (see Section IV-D). During MARL training, the loss function is further augmented with a behavioral constraint term $\mathcal{L}^{\text{BC}}$ to encourage the maintenance of robust multi-hop connectivity (see Section IV-E). Additionally, a knowledge distillation loss $\mathcal{L}^{\text{KD}}$ based on LLMs is designed to promote structured exploration within the MARL paradigm (see Section IV-F). Consequently, the overall objective for each agent $u$ is expressed as
@@ IV-C Information Aggregation for Coordination
To enhance inter-agent cooperation, an information aggregation module is designed for each agent, and its output, together with the agent’s local observation, serves as the input to both the policy and critic networks. Effective cooperation among UAVs in a multi-agent networking environment relies on the exchange of local observations, enabling more informed decision-making and improved coordination. In the networking problem, UAVs are assumed to be able to share their local observations through the established communication network, enabling each UAV to acquire a more comprehensive view of the environment. The shared information at time $t$ is considered as
$\xi(t)=\{$ $\{\boldsymbol{l}_{n}(t)\}_{n\in\boldsymbol{\mathcal{N}}},\{\rho_{i,j}(t)\}_{i,j\in\boldsymbol{\mathcal{N}}},$ (17)
@@
where $\text{concat}(\cdot)$ denotes the concatenation operation, with all vectors flattened before being combined. The local observation of UAV $u$ includes its position $\boldsymbol{l}_{u}$ and the link qualities $\rho_{u,i}$} of all links associated with it. The inclusion of the local observation is essential to ensure that each observation is uniquely associated with a specific UAV, enabling the policy network to map outputs unambiguously to the corresponding decentralized agent. This structure supports stable policy learning and enables agents to have a more comprehensive understanding of the environment, promoting better coordination and decision-making. In practical deployments, this design is feasible as UAVs exchange only compact coordination messages over the multi-hop network. The information aggregation module can be physically separated from the UE service module, thereby reducing interference and improving system reliability.
@@ IV-D2 Relay responsibility
[…]
@@ IV-E Behavioral Constraint for Robustness
[…]
@@
[…]
@@ IV-F LLM Agent and Knowledge Distillation
Given that LLMs possess knowledge and capabilities that are aligned with human preferences, they can be leveraged to guide and enhance the RL training process, thereby reducing unproductive exploration. LLMs do not aim to achieve globally optimal UAV networking solutions, but instead provide high-level strategic guidance that narrows the MARL exploration space. This guidance effectively mitigates the cold-start problem during the initial stages of MARL training, improving learning efficiency and stability. Pretrained LLMs, although primarily trained on natural language data, can understand high-level semantic information about tasks and scenarios. By embedding chain-of-thought (CoT) reasoning into the inference process, LLM-based agents can interpret tasks, extract key features of networking tasks and environmental states, and make informed decisions. Additionally, random initialization in MARL often leads to inefficient exploration in the early stages, heightening the risk of converging to a local optimum. Given the scarcity of feasible network configurations within the vast search space, we propose utilizing LLM-driven decision-making to enhance the MARL training process. Therefore, an LLM agent is designed to offer an initial attempt at tackling the UAV multi-hop networking task, as shown in Figure 3. To ensure that the LLM fully comprehends the networking task and the environmental state for better reasoning, the networking scenario is simplified to preclude the analysis of thousands of precise numerical values, such as location coordinates. And LLM agents generally do not perform at the level of models that have undergone domain-specific training. In this paper, the LLM module is designed primarily to supply RL agents with decisions that align with common sense, thereby aiding their exploration and training processes. Thus, the adoption of an appropriately simplified environment is considered both acceptable and beneficial. Notably, a single UAV can provide communication services over a contiguous region. In line with this regional characteristic of network service delivery, the entire environment is partitioned into grid cells. The grid width is configured to ensure that the UAV positioned at the center of each grid cell can communicate with the UAVs in the adjacent cells and establish connections with the majority of UEs within the grid, i.e. meeting the SNR requirements. Thus, the grid side length is given by $d^{\text{grid}}=\frac{1}{\sqrt{2}}(\max\{d_{u,v}|\rho_{u,v}\geq\rho_{\text{th}}\})$, where $u,v\in\boldsymbol{\mathcal{U}}$ and the SNR $\rho_{u,v}$ is determined by the distance $d_{u,v}$ between UAVs. Consequently, the distribution of UEs is quantified by counting the number of UEs in each grid cell. Moreover, the UAV positions generated by the LLM are constrained to the centers of these grid cells, reducing the complexity of considering connectivity constraints.
@@
Directly mapping current states to UAV positions poses a challenge for the LLM agent [41], as it requires accounting for UE distribution, organization of multi-hop networks and ensuring UAV connectivity constraints for networking. Inspired by the CoT approach [41], the LLM reasoning process for the problem is partitioned into three sequential steps: a) Analyze UE distribution. Identify densely populated areas and determine candidate UAV deployment locations. b) Address connectivity gaps. Evaluate potential UAV connectivity gaps and reorganize UAV placements to ensure robust network connectivity. c) Determine the final UAV deployment based on the analysis of connectivity constraints and UE distribution. This stepwise approach enables the LLM to use more intermediate tokens to sequentially analyze the scenario and requirements, thereby enhancing the overall quality of the response.
To ensure reliable integration of LLM-generated guidance into MARL training, a rule-based verifier is used to validate UAV deployment plans and filter out infeasible outputs. CoT reasoning by the LLM can produce incorrect or inconsistent intermediate steps when generating high-level strategic plans for complex multi-UAV multi-hop networking tasks. The verifier checks that UAV positions are within operational boundaries and physically reachable. It verifies network connectivity by confirming that multi-hop paths remain connected, only a small number of UAVs are isolated, and sufficient UE coverage is maintained according to predefined thresholds. Deployment plans that fail these checks are discarded, preventing infeasible guidance from influencing MARL policy learning. Other methods, such as Retrieval-Augmented Generation [31], Reasoning and Acting [47], or Supervised Fine-Tuning [30], can also improve the reliability of LLM outputs. However, they introduce substantial computational overhead during training. Discarding invalid outputs using a lightweight, rule-based verifier ensures the quality and reliability of LLM-generated guidance while providing an efficient and practical solution for the task.
@@
[…]
@@
[…]
@@
This formulation enables MARL agents to effectively leverage the LLM’s inferred actions as supervisory signals, allowing them to learn the LLM’s decision-making capabilities. In the proposed framework, the LLM participates exclusively during the offline training phase, providing high-level strategic guidance for UAV deployment and multi-hop network coordination. Its outputs are incorporated into MARL policies through a soft target-based distillation loss, ensuring that strategic priors guide the learning process. This offline integration prevents the LLM from participating in real-time UAV control, avoiding additional latency and computational overhead, and making the method feasible for practical network deployments. During online deployment, each UAV executes its MARL policy independently, without LLM inference, ensuring that real-time control is decentralized, incurs no extra computational burden, and remains practical and reproducible in large-scale multi-hop UAV networks. This technique promotes structured exploration and enhances both the training and inference processes of MARL.
IV-G Overall Algorithm and Complexity Analysis
Algorithm 1 Training procedure of MRLMN
Initialize: Policy networks $\pi_{\theta}$, PPO replay buffer $\mathcal{D}$
Parameters: $N_{\text{ep}}$ (episodes), $T$ (episode horizon), $K$ (training epochs), $Q_{\text{LLM}}$ (LLM guidance interval)
Cache: $\mathcal{T}^{\text{LLM}}_{\text{cached}}\leftarrow\emptyset$ $\triangleright$ Cached LLM-suggested UAV deployment
for episode $k=1\dots N_{\text{ep}}$ do
  $s_{0}\sim p_{\text{init}}(s)$ $\triangleright$ Reset environment
  Initialize agent grouping for the episode
  for step $t=1\dots T$ do
   $\mathbf{o}_{t}^{\text{ind}}\leftarrow\text{GetEnvironmentObservation}(s_{t})$
   $\mathbf{o}_{t}\leftarrow\text{AggregateInfo}(\mathbf{o}_{t}^{\text{ind}},\text{grouping})$
   $\triangleright$ Information aggregation
   if $t\bmod Q_{\text{LLM}}==0$ then
     $\mathcal{T}^{\text{LLM}}_{\text{cached}}\leftarrow\text{LLM\_Inference}(\mathbf{o}_{t})$
     $\triangleright$ Update and cache LLM guidance
   end if
   $\mathbf{a}_{t}\sim\pi_{\theta}(\cdot\mid\mathbf{o}_{t})$
   $s_{t+1},\mathbf{r}_{t}\leftarrow\text{Env.Step}(\mathbf{a}_{t})$ $\triangleright$ Environment transition
[…]
   $\triangleright$ Bipartite matching
   $\mathbf{R}_{t}\leftarrow\text{Reward\_Decomp}(\mathbf{r}_{t},\text{grouping})$
   $\triangleright$ Reward Decomposition
   $\mathcal{D}\leftarrow\mathcal{D}\cup\{(\mathbf{o}_{t},\mathbf{a}_{t},\mathbf{R}_{t},\mathbf{o}_{t+1},\sigma^{*})\}$
  end for
  for epoch $1\dots K$ do
   $\max_{\pi_{\theta}}\;\mathcal{L}^{\text{PPO}}-\beta_{1}\mathcal{L}^{\text{KD}}-\beta_{2}\mathcal{L}^{\text{BC}}$
   $\triangleright$ MARL Optimization with knowledge distillation and behavioral constraints
  end for
end for
The complete training procedure of MRLMN is summarized in Algorithm 1. The framework operates in a hybrid manner. UAVs continuously execute decentralized control using MLP-based policy networks $\pi_{\theta}$ under the IPPO paradigm, while the LLM is queried only once every $Q_{\text{LLM}}$ steps to provide high-level deployment guidance. At each step, UAVs first aggregate observations from other UAVs, take actions through the learned policies, and collect transitions into the replay buffer $\mathcal{D}$. When LLM guidance is triggered, the suggested deployment topology is cached using $\mathcal{T}^{\text{LLM}}_{\text{cached}}$ and aligned with the UAVs’ current configuration through bipartite matching, enabling knowledge distillation to shape subsequent policy updates. IPPO optimization is then performed at the end of each episode, incorporating both the distillation objective and the behavioral constraint. From a computational perspective, each UAV is equipped with an MLP consisting of $L$ hidden layers of width $H$. Thus, with $U$ UAVs executing their policies in parallel, the per-step inference cost scales as $\mathcal{O}(U\cdot L\cdot H^{2})$. During training, the same order applies to forward and backward passes of both the policy and critic networks. LLM reasoning is executed once every $Q_{\text{LLM}}$ steps, with a single inference cost $\mathcal{C}_{\text{LLM}}$ contributing an amortized $\mathcal{C}_{\text{LLM}}/Q_{\text{LLM}}$ per step. The bipartite matching needed for distillation incurs $\mathcal{O}(U^{3})$ complexity, which is lower than the neural network computations and is invoked once per step. Additionally, the overhead from information aggregation, reward decomposition, and behavioral constraints is minimal. Consequently, MLP-based networks dominate inference complexity, whereas during training, both the neural networks and the amortized LLM guidance jointly determine the overall computational cost.
@@ V Performance Evaluation
This section evaluates the proposed MRLMN framework in multi-UAV multi-hop networking scenarios under diverse conditions, including varying environment scales, UAV swarm sizes, baseline comparisons, training dynamics, and ablation studies. The experiments demonstrate the framework’s stability, scalability, adaptability, and coordination effectiveness, while also examining the impact of parameter sharing within agent groups on training efficiency and policy quality. The results highlight the overall superior performance of MRLMN across these settings.
@@ V-A1 Environment setup
To assess the performance of our proposed method, we employ a simulation environment that spans an area of approximately 3.5 km $\times$ 3.5 km, within which around 150 UEs and 18 UAVs are deployed. The distribution of UEs follows either a uniform pattern or a two-dimensional Gaussian mixture with multiple centers, and their motion follows a Brownian process at a constant velocity. The BSs are strategically positioned at three of the four corners of the area to preclude direct connections with the UEs, thereby simulating communication conditions typically encountered in disaster scenarios. The UAV speed $\omega$ is limited to a displacement of 30 meters within each time slot.